OpenStudy (anonymous):

7. Solve the equation 5/4x-4 - 6/x+1 = -1 Help Please!!!

OpenStudy (paki):

please write the equation again... i am not getting it... pl

OpenStudy (anonymous):

(5/4x-4) - (6/x+1) = -1

OpenStudy (anonymous):

wait i gtg srry

OpenStudy (paki):

ok

OpenStudy (paki):

just multiply both sides of the equation by (4x-4)(x++1), to make it simple... then solve..

OpenStudy (anonymous):

help

OpenStudy (mathmale):

paki has identified the lowest common denominator and has asked you to multiply all terms of the given equation by it. This is good advice. Please follow it.

OpenStudy (anonymous):

I didn't get what paki meant

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

ok so 5 - 6/3x+3 = -4x-4 ???

OpenStudy (anonymous):

hello?

OpenStudy (mathmale):

If you actually do multiply the original equation by the LCD identified by paki, you would no longer have fractions. Please go back and try that again. Write down the original equation. Then multiply each term of that equation by the LCD.

OpenStudy (anonymous):

wait the LCD is 4x+4 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so -7 = -4x-4 ???

OpenStudy (anonymous):

-3=-4x

OpenStudy (anonymous):

3/4 = x ???

OpenStudy (anonymous):

mathmale help?

OpenStudy (paki):

@ SwagBag see... (5/4x-4)-(6/x+1)=-1 multiply both sides of the equation, to make it simple... (4x-4)(x+1)*(5/4x-4)-(6/x+1)(4x-4)(x+1)=-1(4x-4)(x+1) 5(x+1)-(4x-4)6=-((4x-4)(x+1) 5x+1-24x+24=-4x^2+4x-4x+4 +5x-24x+4x^2=+4-24-1 19x+4x^2=-19 4x^2+19x-19=0 now factorize or use quadratic formula to solve...

OpenStudy (mathmale):

@swagbag : Would you please consider getting and sharing a screen capture of this problem ... 5/4x-4 - 6/x+1 = -1 Without parentheses, or even with the parentheses you provided later, this problem statement is subject to misinterpretation. If you post a screen shot, or if you use your cell phone camera to photograph the problem, and share the image with me, I will help you again.